OpenStudy (anonymous):

The amount of money an electrician charges when visiting a home to make a repair is shown by the linear equation c = 110 + 85h, where c is the amount the electrician charges and h is the number of hours spent at the home. How much does the electrician charge when h = 3.5?

OpenStudy (bibby):

what is 85 h?

OpenStudy (anonymous):

297.5

OpenStudy (bibby):

what is c?

OpenStudy (anonymous):

407.5 ?

OpenStudy (bibby):

c = 110 + 85h 85h=297.5 c=110+297.5 c=407.5 these are repetitive and you know them, why do you want your hand held through them?

OpenStudy (anonymous):

Eh Sorry I did my 4th question and got it wrong so just wanted to make sure this time

OpenStudy (bibby):

sorry for making you apologize for just checking your work

OpenStudy (anonymous):

pf it's fine bibsu

OpenStudy (bibby):

many thanks this time kay kay cee

OpenStudy (anonymous):

Omfg " kay kay cee " that's gr8 m8 many welc 2 u bibsu

OpenStudy (bibby):

I want to be able to talk to you in single letters ok?

OpenStudy (bibby):

gmyw2ub = gr8 m8, you're welcome 2 you bibsu

OpenStudy (bibby):

that isn't a single letter I'm an idiot

OpenStudy (bibby):

ignore me

OpenStudy (anonymous):

Omg I'm laughing w o w

OpenStudy (bibby):

you make me blush but I can't blush bc brown so ya know on the inside I get all tingly for a second you make me feel weird im going to ask my mom what this is all about

OpenStudy (anonymous):

oh my why

OpenStudy (bibby):

no reason nothing ever happens for a why they just happen

OpenStudy (bibby):

you'll understand when you're older

OpenStudy (anonymous):

I know but eh I have a girlfriend just to let you know

OpenStudy (bibby):

rofl wut

OpenStudy (bibby):

thank you random internet dude who I can't touch

OpenStudy (bibby):

I was romantically interested in you :(

OpenStudy (anonymous):

Oh my I'm not a dude laughs

OpenStudy (bibby):

I KNOW THAT YOU IDIOT

OpenStudy (bibby):

JESUS CHRIST MY POINT IS I CAN'T TOUCH YOU AND THUS HAVE NO ROMANTIC INTEREST IN YOU I DON'T EVEN KNOW WHAT YOU LOOK LIKE stop making me yell

OpenStudy (anonymous):

o gosh okay sorry m8

OpenStudy (anonymous):

u wanna see a picture of me laughs jknonever

OpenStudy (bibby):

I do u wanna see my collection of mattyb pictures?

OpenStudy (anonymous):

mfg why

OpenStudy (bibby):

don't show this hot porn to any1

OpenStudy (bibby):

I hope we get in trouble for this

OpenStudy (anonymous):

O my that really rustles my jimmies

OpenStudy (bibby):

I'm reporting myself

OpenStudy (anonymous):

I just choked on my gum bcuz of that

OpenStudy (bibby):

this is my hand doing things with a cat

OpenStudy (bibby):

it looks chubby but I assure you I am a normal boy

OpenStudy (anonymous):

o gog h e l p can't breath properluy

OpenStudy (bibby):

do you want to keep going through my desktop images?

OpenStudy (bibby):

is it weird for a 21 year old man to have this picture

OpenStudy (anonymous):

jfc that's aAMAZING w 0 w

OpenStudy (bibby):

i thot u were gay please educate me on tumblr se*uality

OpenStudy (anonymous):

He looks like a little transgender babs - also, everyone on tumblr is openly pansexual/bi/married to their pet hedgehog it's law

OpenStudy (bibby):

I'm a hedgehog kin myself, pleasure 2 meat u

OpenStudy (bibby):

do you wanna go halfsies on this one dude?

OpenStudy (bibby):

ugh

OpenStudy (bibby):

am I distracting you from homework

OpenStudy (anonymous):

yes i'm free and up 2 that a n y t i m e

OpenStudy (anonymous):

oh yes, yes u r

OpenStudy (bibby):

let's do your homework I promise it'll be just as entertaining

OpenStudy (anonymous):

wow ok , he last question i have for the night i just need checked

OpenStudy (bibby):

leaked footage of me in real life

OpenStudy (bibby):

we doing this or what kkc

OpenStudy (anonymous):

o wow ur fabulous

OpenStudy (bibby):

thank....

OpenStudy (anonymous):

welc

OpenStudy (bibby):

post the freaking question wow

OpenStudy (bibby):

oh nvm

OpenStudy (anonymous):

omg I'm in the middle of posting it shush

OpenStudy (bibby):

dude I said sorry

OpenStudy (bibby):

in my own way

OpenStudy (anonymous):

ik ik okay im posting it now bibsu

OpenStudy (anonymous):

can you guys just give the answer oh my god